How many liters of a 40% solution and an 16% solution must be mixed to obtain 20 liters of a 22% solution?
Answer by stanbon(48558)

How many liters of a 40% solution and 16% solution must be mixed to obtain 20 liters of a 22% solution?
-------------------
Equation:
0.40x + 0.16(20-x) = 0.22(20)
---------------------------------
Multiply thru by 100 to get:
40x + 16*20 - 16x = 22*20
24x = 6*20
x = 5 liters (amount of 40% solution needed in the mixture)
---------
20-x = 20-5 = 15 liters (amount of 16% solution needed in the mixture)
===========================
Cheers,
Stan H.